OpenStudy (anonymous):

f'(x) =-4sin(x) + 7sec^2(x) find f'(4pi/3)

OpenStudy (tkhunny):

Have you considered substituting \(x = 4\pi/3\) into your equation for f'(x)?

OpenStudy (anonymous):

yes i already found the prime but i dont know how it would work out since there is sin and cos @tkhunny

OpenStudy (anonymous):

You are given what f'(x) is and you're asked to find f'(4pi/3). So just plug in x = 4pi/3, as @tkhunny suggested, in to f'(x) and evaluate... @foreveryounging97

OpenStudy (tkhunny):

\(\sin(4\pi/3) = -\sin(\pi/3) = -\sqrt{3}/2\)?

OpenStudy (anonymous):

18/7

OpenStudy (tkhunny):

\(\sec(4\pi/3) = 1/\cos(4\pi/3) = -1/\cos(\pi/3) = -1/(1/2) = -2\) Perhaps you should back up a step and brush up on your trigonometry?